Use the equation -6x+9y=18 to identify the slope and the y-intercept of the graph.

Answer by sudhanshu_kmr(1152)


-6x + 9y = 18
=> 9 y = 6x + 18
=> y = 6/9 x + 18/9
=> y = 2/3 x + 2
so, slope = 2/3 , y-intercept = 2